# 2d Conformal Field Theories on Magic Triangle

### v1

@@ 5 Magic triangle at arbitrary positive integer levels
Denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$. For the 16 elements of the magic triangle that are semisimple Lie algebras $\mathfrak{g}$, the corresponding RCFT with arbitrary positive integer level $k$ are just their WZW models $(\mathfrak{g})_{k}$, except for ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$ for which the levels must be rescaled by $3$, 88, and $2$, respectively. For the remaining 14 elements, a case-by-case analysis is required. By carefully examining the central charges and conformal weights, we arrive at the following conclusions:
@@
  • $\mathcal{T}^{(1/3,4)}_{k}$ is the effective minimal model $\mathrm{Vir}_{3k+2,3}^{\rm eff}$. This has been realized in Lee et al. (2024b).
  • $\mathcal{T}^{(1/2,3)}_{k}$ is the compact boson $U(1)_{3k}$ with $c=1$.
@@
  • $\mathcal{T}^{(1,2)}_{k}$ is a non-diagonal modular invariant of $U(1)_{2k}\times U(1)_{6k}$. The vacuum character is given by
    $\chi^{U(1)_{2k}}_{0}\chi^{U(1)_{6k}}_{0}+\chi^{U(1)_{2k}}_{\frac{k}{2}}\chi^{U(1)_{6k}}_{\frac{3k}{2}}.$ (115)
@@
  • $\mathcal{T}^{(4,5)}_{k}$, $\mathcal{T}^{(4,3)}_{k}$, $\mathcal{T}^{(4,2)}_{k}$ exist up to $k=5,3,2$ respectively. These theories are equivalent to the minimal $\mathcal{W}$-algebras of E-type with negative integer levels, which have recently been studied in Sun (2024); Arakawa et al. (2024); Kac and Wakimoto (2025). Other $\mathcal{T}^{(4,\nu)}_{k}$ theories for $k>1$ should not exist.
@@ 6 Concluding remarks
In this work, we have extended the magic triangle and identified the two-dimensional conformal field theories associated with its elements at arbitrary positive integer levels. In particular, for level-one theories we uncover a number of remarkable and highly structured properties, including modular linear differential equations, the magic coset construction, five atomic theories, magic bilinear relations, and matrix-valued modularity. For level-two theories, we focus on the subexceptional series and found they are all supersymmetric and their fermionic characters satisfy a uniform fermionic MLDE. We also find many interesting coset constructions involving the Cvitanović–Deligne, sub and Severi exceptional series at level two.
@@
We have not explored much of the dimension of irreducible representations for the magic triangle in this work. The magic coset we have uncovered suggest that the representations of all algebras appearing in the magic triangle are intricately entangled. It would be interesting to find some simple universal formulas for the dimension and dual Coxeter number of these representations, besides the adjoint representation.
Finally, although our discussion has focused on WZW models at positive integer levels, theories at fractional and negative levels are also of significant interest. For instance, it is well known that the Cvitanović–Deligne exceptional series at level $h^{\vee}/6-1$ gives rise to a celebrated family of quasi-lisse vertex operator algebras, whose vacuum characters satisfy a one-parameter family of second-order MLDEs Beem and Rastelli (2018); Arakawa and Kawasetsu (2018). It is intriguing to consider whether this phenomenon can be generalized to other exceptional series, or even to the entire magic triangle.
@@ Acknowledgements
We thank Chongying Dong, Heeyeon Kim, Ching Hung Lam, Jaewon Song, Haowu Wang, Xin Wang, Yinan Wang and Don Zagier for discussion. KL is supported in part by the BIMSA start-up fund and the Beijing Natural Science Foundation International Scientist Project No.IS25024. KS is supported by NSF No.1250010384.
@@ Appendix A Explicit data of the extended magic triangle
We collect the dual Coxeter number of dimension of each element of the extended magic triangle in Table 5. We collect the dimensions of $R_{1},R_{2},R_{3}$ for all theories at level one in Table 6. We collect the central charge, nonzero conformal weights and degeneracies for all theories at level one in Table 7. Due to the symmetry between $\mu$ and $\nu$, we only show the $\mu\leq\nu$ part of the triangle.
@@ Appendix B S-matrices
In this section, we collect the $S$-matrices for the two characters $\chi_{0},\chi_{1}$ of the Mathur–Mukhi–Sen series, i.e. the extended Cvitanović–Deligne exceptional series at level one, reviewed in Section 2.3. The $2\times 2$ $S$-matrix $S_{(\nu)}$ for the two characters of $\mathcal{T}^{(\nu,5)}$ which leaves the partition function (15) invariant satisfies $S_{(\nu)}=S_{(1/\nu)}.$ The $S$-matrices for $\nu=1/4,1/3,1/2,2/3,1$ of ${\rm Vir}_{5,2}^{\rm eff},(A_{1})_{1},(A_{2})_{1},(G_{2})_{1},(D_{4})_{1}$ theories are, respectively,
$S_{(4)}=\frac{\varphi^{1/2}}{5^{1/4}}\left(\begin{array}[]{cc}1&\varphi^{-1}\\ \varphi^{-1}&-1\end{array}\right),\quad S_{(3)}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cr}1&1\\ 1&-1\end{array}\right),\quad S_{(2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{cr}1&2\\ 1&-1\end{array}\right),$ (122)
$S_{(3/2)}=\frac{1}{5^{1/4}\varphi^{1/2}}\left(\begin{array}[]{cc}1&\varphi\\ \varphi&-1\end{array}\right),\quad S_{(1)}=\frac{1}{2}\left(\begin{array}[]{cr}1&3\\ 1&-1\end{array}\right),$ (127)
where $\varphi=(\sqrt{5}+1)/2$ is the Golden ratio. The partition function $Z=|\chi_{0}|^{2}+\mathcal{N}|\chi_{1}|^{2}$ is invariant under these $S$ matrices. Note that
@@
For the cases with degeneracy, i.e. $\nu=1/2,1,2$, we can have extended $S$-matrices for $\nu=1/2,1$ as
$\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}1&1&1\\ 1&w&w^{2}\\ 1&w^{2}&w\end{array}\right),\quad\frac{1}{2}\left(\begin{array}[]{crrr}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right),$ (140)
@@ Appendix C Spurious solutions of MLDE
For each $\mu$, besides the two characters of $\mathcal{T}^{(5,\mu)}$, there are two spurious solutions of the fourth order MLDE (41) with Fourier coefficients with unbounded denominators. For example, for $\mu=1/4$, the two genuine solutions are the Lee–Yang characters $\chi_{0},\chi_{1/5}$, while the two spurious solutions are
@@
Although these two solutions are not characters themselves, we observe that they are closely related to the characters of the $(E_{7+1/2})_{1}$ theory:
@@
These relations are easy to prove since $\Delta_{4}$ annihilates both sides of the equations. If we define $\chi_{\frac{1}{30}}=R_{1}f_{\frac{1}{30}}$ and $\chi_{\frac{5}{6}}=R_{3}f_{\frac{5}{6}}$, by properly choosing the prefactors such that $R_{3}/R_{1}=3/5$, we find a bilinear relation
@@
similar to the magic bilinear relation (46). This choice also makes the relation (47) hold. For other $\mu\neq 5$, we obtain similar relations.
The level one $E_{8}$ theory has only one character $\chi_{0}=j^{1/3}(\tau)$ where $j(\tau)$ is the Klein $j$-function. The vacuum character satisfies the first order MLDE,
@@
The second order MLDE (10) for $E_{8}$ theory gets factorized as
@@
For $\mu,\nu=5$, the 4th order MLDE (41) becomes
$\left[D^{4}-\frac{1}{6}E_{4}D^{2}+\frac{1}{9}E_{6}D-\frac{1}{36}E_{4}^{2}\right]\chi=0.$ (149)
It factorizes to $D^{2}(D^{2}-\frac{1}{6}E_{4})]\chi=0$. Beside $\chi_{0}=j^{1/3}$ the vacuum character of $(E_{8})_{1}$, this MLDE has two more spurious solutions
@@
These are not characters, but satisfy the following simple relations
@@
We find a bilinear relation

### v2

@@ 5 Magic triangle at arbitrary positive integer levels
We denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$. For the 16 elements of the magic triangle that are semisimple Lie algebras $\mathfrak{g}$, the corresponding RCFT with arbitrary positive integer level $k$ are just their WZW models $(\mathfrak{g})_{k}$, except for ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$ for which the levels must be rescaled by factors $3$, 88, and $2$, respectively. For the remaining 14 elements, a case-by-case analysis is required. By carefully examining the central charges and conformal weights, we arrive at the following conclusions:
@@
  • $\mathcal{T}^{(1/3,4)}_{k}$ is the effective minimal model $\mathrm{Vir}_{3k+2,3}^{\rm eff}$. This was realized in Lee et al. (2024b).
  • $\mathcal{T}^{(1/2,3)}_{k}$ is the compact boson $(U_{1})_{3k}$ with $c=1$. The central charge is due to $d_{\mathfrak{g}}=1$ and $h^{\vee}=0$.
@@
  • $\mathcal{T}^{(1,2)}_{k}$ is a non-diagonal modular invariant of $(U_{1})_{2k}\times(U_{1})_{6k}$ with $c=2$. The central charge is due to $d_{\mathfrak{g}}=2$ and $h^{\vee}=0$. The vacuum character is given by
    $\chi^{(U_{1})_{2k}}_{0}\chi^{(U_{1})_{6k}}_{0}+\chi^{(U_{1})_{2k}}_{\frac{k}{2}}\chi^{(U_{1})_{6k}}_{\frac{3k}{2}}.$ (115)
@@
  • The theories $\mathcal{T}^{(4,5)}_{k}$, $\mathcal{T}^{(4,3)}_{k}$, and $\mathcal{T}^{(4,2)}_{k}$ exist up to $k=5,3,2$ respectively. These theories are equivalent to the minimal $\mathcal{W}$-algebras of E-type with negative integer levels, which have recently been studied in Arakawa and Moreau (2018); Sun (2024); Arakawa et al. (2024); Kac and Wakimoto (2025). Other $\mathcal{T}^{(4,\nu)}_{k}$ theories for $k>1$ are not expected to exist.
@@ 6 Concluding remarks
In this work, we have extended the magic triangle and identified the two-dimensional conformal field theories associated with its elements at arbitrary positive integer levels. In particular, for level-one theories we uncover a number of remarkable properties, including modular linear differential equations, the magic coset, five atomic theories, magic bilinear relations, the dimensions and degeneracies of primary operators, and matrix-valued modularity including $S$-matrices. For level-two theories, we focus on the subexceptional series and found they are all supersymmetric and their fermionic characters satisfy a uniform fermionic MLDE. We also find many interesting coset constructions involving the Cvitanović–Deligne, sub and Severi exceptional series at level two.
@@
We have not explored much of the dimension of irreducible representations for the magic triangle in this work. The magic coset we have uncovered suggest that the representations of all algebras appearing in the magic triangle are intricately interconnected. It would therefore be interesting to find some simple universal formulas for the dimensions of these representations, besides the adjoint representation and some simpler representations appearing as those for non-vacuum primary operators.
Finally, although our discussion has focused on WZW models at positive integer levels, theories at fractional and negative levels are also of considerable interest. For instance, it is well known that the Cvitanović–Deligne exceptional series at level $h^{\vee}/6-1$ gives rise to a celebrated family of quasi-lisse vertex operator algebras, whose vacuum characters satisfy a one-parameter family of second-order MLDEs Beem and Rastelli (2018); Arakawa and Kawasetsu (2018). It is intriguing to consider whether this phenomenon can be generalized to other exceptional series, or even to the entire magic triangle.
@@ Acknowledgements
We thank Chongying Dong, Heeyeon Kim, Ching Hung Lam, Jaewon Song, Haowu Wang, Xin Wang, Yinan Wang and Don Zagier for discussion. KL is supported in part by the BIMSA start-up fund and the Beijing Natural Science Foundation International Scientist Project No.IS25024. KS is supported by China NSF grant No.1250010384.
@@ Appendix A Explicit data of the extended magic triangle
We collect the dual Coxeter number of dimension of each element of the extended magic triangle in Table 5. We collect the dimensions of $R_{1},R_{2},R_{3}$ for all level-one theories in Table 6. The central charge, nonzero conformal weights and degeneracies for all level-one theories are summarized in Table 7. Owing to the symmetry exchanging $\mu$ and $\nu$, we display the $\mu\leq\nu$ portion of the triangle.
@@ Appendix B S-matrices
In this section, we collect the $S$-matrices for the two characters $\chi_{0},\chi_{1}$ of the Mathur–Mukhi–Sen series, namely the extended Cvitanović–Deligne exceptional series at level one, as reviewed in Section 2.3. The $2\times 2$ $S$-matrix $S_{(\nu)}$ for the two characters of $\mathcal{T}^{(\nu,5)}$, which leaves the partition function (15) invariant, satisfies $S_{(\nu)}=S_{(1/\nu)}.$ The $S$-matrices for $\nu=1/4,1/3,1/2,2/3,1$, corresponding respectively to the theories ${\rm Vir}_{5,2}^{\rm eff}$, $(A_{1})_{1}$, $(A_{2})_{1}$, $(G_{2})_{1}$, $(D_{4})_{1}$, are given by
$S_{(1/4)}=\frac{\varphi^{1/2}}{5^{1/4}}\left(\begin{array}[]{cc}1&\varphi^{-1}\\ \varphi^{-1}&-1\end{array}\right),\quad S_{(1/3)}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cr}1&1\\ 1&-1\end{array}\right),\quad S_{(1/2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{cr}1&2\\ 1&-1\end{array}\right),$ (122)
$S_{(2/3)}=\frac{1}{5^{1/4}\varphi^{1/2}}\left(\begin{array}[]{cc}1&\varphi\\ \varphi&-1\end{array}\right),\quad S_{(1)}=\frac{1}{2}\left(\begin{array}[]{cr}1&3\\ 1&-1\end{array}\right),$ (127)
where $\varphi=(\sqrt{5}+1)/2$ is the golden ratio. The modular invariant partition function is $Z=|\chi_{0}|^{2}+\mathcal{N}|\chi_{1}|^{2}$ and $\mathcal{N}$ is the degeneracy of the non-vacuum primary fields. Note that
@@
For the cases with degeneracy, namely $\nu=1/2,1,2$, we can have extended $S$-matrices, which are given by
$S_{(1/2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}1&1&1\\ 1&w&w^{2}\\ 1&w^{2}&w\end{array}\right),\quad\ S_{(1)}=\frac{1}{2}\left(\begin{array}[]{crrr}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right),$ (140)
@@ Appendix C Spurious solutions of MLDE
For each $\mu$, in addition to the two genuine characters of $\mathcal{T}^{(5,\mu)}$, there exist two spurious solutions of the fourth order MLDE (41) whose q-expansion have coefficients with unbounded denominators. As an illustrative example, for $\mu=1/4$, the two genuine solutions are the Lee–Yang characters $\chi_{0},\chi_{1/5}$, while the two spurious solutions are given by
@@
Although these two solutions are not themselves RCFT characters, we observe that they are closely related to the characters of the $(E_{7+1/2})_{1}$ theory via the relations
@@
Here $\Delta_{\mu}$ is the second order differential operator defined in (14). These relations are straightforward to prove, since $\Delta_{4}$ annihilates both sides of the equations. If we define $\chi_{\frac{1}{30}}=R_{1}f_{\frac{1}{30}}$ and $\chi_{\frac{5}{6}}=R_{3}f_{\frac{5}{6}}$, and choose the prefactors such that $R_{3}/R_{1}=3/5$, we obtain the bilinear relation
@@
similar to the magic bilinear relation (46). This choice also ensures the relation (47) holds. For other $\mu\neq 5$, similar relations can be derived.
The level one $(E_{8})_{1}$ theory has a single character $\chi_{0}=j^{1/3}(\tau)$ where $j(\tau)$ is the Klein $j$-invarant. The vacuum character satisfies the first order meromorphic MLDE,
@@
Accordinngly, the second-order MLDE (10) for the $E_{8}$ theory factorizes as
@@
For $\mu,\nu=5$, the fourth order MLDE (41) becomes
$\left[D^{4}-\frac{1}{6}E_{4}D^{2}+\frac{1}{9}E_{6}D-\frac{1}{36}E_{4}^{2}\right]\chi=\left[D^{2}(D^{2}-\frac{1}{6}E_{4})\right]\chi=0.$ (149)
Besides $\chi_{0}=j^{1/3}$ the vacuum character of $(E_{8})_{1}$, this MLDE has two more spurious solutions
@@
These solutions are not characters, but satisfy the following simple differential equations
@@
We find a bilinear relation for the three solutions